According to a survey on US population, the probability that a man will live seventy-five years is 0.62 and that he will live up to eighty-five years is 0.023. If Mr. Clooney is seventy-five years old, what is the probability he will survive to eighty-five years?
Express the probability as a percentage. Include in your final answer the calculations necessary to arrive at the probability

Answer by Theo(13342)

the probability someone will live to 75 years of age is .62
the probability someone will live to 85 years of age is .023

these probabilities are based on the total population.

clooney made it to 75.

what is the probability that he will live to be 85 given that he already made it to 75.

the probability should be .023 / .62 = .0370967742

why is this?

assume there are 1000 men who were born at the same time.
.62 * 1000 = 620 of them will make it to 75.
.023 * 1000 = 23 of them will make it to 85.

the proportion of those who will make it to 85 assuming that they already made it to 75 will be 23 / 620 = .0370967742

multiply that by 100 and you get the percent.

the percent is therefore equal to 3.70967762% which can be rounded to the number of decimal places you think is appropriate or is required.

for example, rounding to one decimal place gets you 3.7%.